OpenStudy (anonymous):

What is a recursive rule for the sequence 9,14,24,44,84?

jimthompson5910 (jim_thompson5910):

You're starting with 9. Double it to get 9*2 = 18. Then subtract 4 to get 18-4 = 14 Then take 14 and double it: 14*2 = 28, then subtract 4: 28 - 4 = 24 Then take 24 and double it: 24*2 = 48, then subtract 4: 48 - 4 = 44 So the next term is found by doubling the previous term and subtracting 4 So the recursive rule is a[n] = 2*a[n-1] - 4

OpenStudy (anonymous):

Thanks!! =)

jimthompson5910 (jim_thompson5910):

sure thing :)